# MAViS: A Multi-Agent Framework for Long-Sequence Video Storytelling

### v3

@@
Abstract.
Despite recent advances, long-sequence video generation frameworks still suffer from significant limitations: poor assistive capability, suboptimal visual quality, and limited expressiveness. To mitigate these limitations, we propose MAViS, an end-to-end multi-agent collaborative framework for long-sequence video storytelling. MAViS orchestrates specialized agents across multiple stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.In each stage, agents operate under the 3E Principle—Explore, Examine, and Enhance—to ensure the completeness of intermediate outputs. Considering the capability limitations of current generative models, we propose the Script Writing Guidelines to optimize compatibility between scripts and generative tools. Experimental results demonstrate that MAViS achieves state-of-the-art performance in assistive capability, visual quality, and video expressiveness. Its modular framework further enables scalability with diverse generative models and tools. With just a brief user prompt, MAViS is capable of producing high-quality, expressive long-sequence video storytelling, significantly improving the assistive capability of enriching inspirations and creativity for users. To the best of our knowledge, MAViS is the only framework that provides multimodal design output – videos with narratives and background music.
$\dagger$ Corresponding author and project lead
1. Introduction
In recent years, video generative models (OpenAI, 2024; MiniMax, 2024; GenmoTeam, 2024) have made substantial strides in text-to-video (T2V) and image-to-video (I2V) generation. While these methods demonstrate promising capabilities for shot-level creation, the generation of long videos remains a formidable challenge. Current scale of compute and data is insufficient to train a single end-to-end model capable of producing high-quality, minute-long videos.
Table 1. Properties of MAViS vs Other Methods: We summarize seven metrics for long-sequence video storytelling works.”Agentic”: whether the system employs a multi-agent architecture; ”Script Writing”: ability to automatically generate scripts from user prompt; ”Shot Designing.”: support for structured shot designing and controllable narrative flow; ”Dynamic”: whether the generated video exhibits dynamic features; ”ID Consistency”: consistency in character identity across multiple shots; ”Audio”: whether the output includes synchronized audio; ”Automated”: whether the pipeline can operate without the presence of user input in script writing, script planning, or model fine-tuning.
Methods Agentic Script Writing Shot Designing Dynamic ID Consistency Audio Automated
Mora (Yuan et al., 2024) $\times$ $\times$ $\times$ $\times$
AesopAgent (Wang et al., 2024c) $\times$ $\times$ $\times$
MovieDreamer (Zhao et al., 2024) $\times$ $\times$ $\times$ $\times$ $\times$
DreamFactory (Xie et al., 2024) $\times$ $\times$ $\times$
VGoT (Zheng et al., 2025) $\times$ $\times$ $\times$
LCT (Guo et al., 2025) $\times$ $\times$ $\times$ $\times$ $\times$
MovieAgent (Wu et al., 2025) $\times$ $\times$
@@ 1. Introduction
To overcome this limitation, prior works have explored extending existing video clips (KlingAI, 2024; Henschel et al., 2024; Ge et al., 2022), or composing sequences of keyframes (Liu et al., 2024; Yang et al., 2024a; Rahman et al., 2023; Wang et al., 2024c), or integrating storytelling and video generation (Yin et al., 2023b; Zheng et al., 2025). However, these approaches often suffer from repetitive motion, lack of narrative cohesion, or result in static image series with limited visual expressions. To better align narrative planning with video animation, MovieDreamer (Zhao et al., 2024) adopts an auto-regressive approach to generate key frames from input scripts, which are then expanded into videos via I2V models, while Dreamfactory (Xie et al., 2024) and MovieAgent (Wu et al., 2025) propose a hierarchical script–scene–shot pipeline. However, as shown in Table 1, these methods require substantial manual effort from users for script writing and training LoRA models to maintain character identity consistency, which significantly limits their assistive capability and hinders large-scale deployment.
@@
In summary, current long-sequence video storytelling methods face three major limitations: 1) poor assistive capability, not supporting script writing, shot designing, and LoRA training to maintain character ID consistency; 2) suboptimal visual quality, manifested in visual distortions, unnatural actions, and inaccurate object proportions; and 3) limited expressiveness, often resulting in repetitive actions and poorly structured shot compositions.
@@
To address this, we introduce MAViS, a multi-agent framework for end-to-end long-sequence video generation. MAViS encompasses various stages including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation. Given a brief user prompt, the script writing stage first generates a structured script with character definitions and shot outlines. The shot designing stage defines a series of shot-level elements to guide subsequent content generation, while character modeling ensures identity consistency via LoRA-trained appearances. During the keyframe generation stage, T2I models are employed to create the initial static frames for each shot, i.e. the keyframes, which are then expanded into full video sequences in the video animation stage through I2V models, maintaining visual coherence and narrative fluency. Finally, voice-overs and background music are automatically added to each shot during the voice generation stage, resulting in a cohesive long-sequence video.
@@
Crucially, all stages operate under the 3E Principle: - Explore: Initial generation of candidate outputs; - Examine: Reviewing the quality and completeness of the output; - Enhance: Iterative refinement and optimization. After multiple iterations, each stage passes the most complete version of its output to the next, enabling gradual optimization throughout the end-to-end pipeline.
@@
To mitigate the limitations of current generative models in maintaining background consistency, executing complex actions, and rendering fine-grained details, we propose the Script Writing Guidelines aimed at enhancing the compatibility between scripts and generative models, ultimately improving the expressiveness of the video storytelling. With just a brief user prompt, MAViS efficiently generates high-quality and complete long-sequence video storytelling, significantly improving the assistive capability of enriching inspirations and creativity for users.Furthermore, its modular architecture ensures scalability with various generative tools and models.
@@
$\bullet$ We propose MAViS, an end-to-end multi-agent framework that encompasses various stages including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation — forming a complete and scalable workflow for long video storytelling.
@@
$\bullet$ Experiments demonstrate that MAViS achieves state-of-the-art performance and win the most user preferences in terms of expressiveness and visual quality. With just a brief user prompt, MAViS efficiently generates high-quality and complete long-sequence video storytelling. To the best of our knowledge, MAViS is the only framework that provides multimodal design output.
2. Related Works
2.1. Image Generation
Recent advances in image generation (Li et al., 2019; Xu et al., 2018; Zhang et al., 2017; Ramesh et al., 2021) have been propelled by deep learning techniques.To fulfill the task of visual storytelling, some works (Liu et al., 2024; Yang et al., 2024a; Rahman et al., 2023) generate image sequences from text and image-caption pairs. To ensure subject consistency (Gal et al., 2022), DreamBooth (Ruiz et al., 2023) adopts subject-specific fine-tuning, while GPT-Image-1 (OpenAI, 2025) introduces context-aware generation for identity preservation. In this work, MAViS leverages off-the-shelf models to generate keyframes for each shot, which serves as the basis for subsequent video animation.
Figure 2. Illustration of MAViS framework. See additional presentation is in the supplementary video.
2.2. Video Generation
@@
Long video generation (He et al., 2022; Ge et al., 2022; Wang et al., 2024e) aims to produce longer videos than previous text-to-video (T2V) (ModelScopeTeam, 2023; Wang et al., 2024b) and image-to-video (I2V) (Girdhar et al., 2023; Xing et al., 2025; Zhang et al., 2023) models. Auto-regressive approaches (Weng et al., 2024; Henschel et al., 2024) extend videos by conditioning on earlier frames, yet often result in repetitive motions with limited narrative development. Attempts to directly integrate storytelling and video generation (Yin et al., 2023b; Zheng et al., 2025) suffer from poor visual quality and unstructured video composition, leading to subpar viewer experiences. Meanwhile, commercial models have achieved notable success in both T2V (GenmoTeam, 2024; Kong et al., 2024; OpenAI, 2024) and I2V generation (Google, 2024; Luma, 2024; MiniMax, 2024; Ang and Baole, 2025), and some  (KlingAI, 2024; Runway, 2024a, b) support temporal extension. However, the duration of generated videos remains constrained. In this work, MAViS uses cutting-edge video generation models to synthesize high-quality video clips and assembles them into long-sequence video storytelling.
2.3. AI Agents
@@
AI agents, empowered by large language models (Brown et al., 2020; Touvron et al., 2023a, b), multimodal architectures, and reinforcement learning, have shown broad applicability in reasoning, planning, web interaction, and visual generation. General-purpose LLMs underpin agent frameworks (Yang et al., 2023b; Yin et al., 2023a; Li et al., 2023; Hong et al., 2023; Park et al., 2023), while multimodal models (Achiam et al., 2023; Anil et al., 2023) extend agent capabilities across modalities (Yang et al., 2023a; Zhang and Zhang, 2024; Zheng et al., 2024; Wang et al., 2024a; Liu et al., 2025; Wang et al., 2024d; Hang et al., 2024; Liao et al., 2025). In video generation, systems (Yang et al., 2024b; Yuan et al., 2024; Wang et al., 2024c) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines like Dreamfactory (Xie et al., 2024) and MovieAgent (Wu et al., 2025) introduce script-to-shot workflows but still rely on manual scripting and LoRA tuning. In contrast, our MAViS leverages multi-agent collaboration to enable long-sequence video storytelling.
3. MAViS
@@
We propose MAViS, a multi-agent end-to-end framework for long-sequence video storytelling that consists of various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.
@@
Given a user prompt $P$, MAViS aims to generate a long-sequence video storytelling that comprises $N$ shots:
(1) $\mathcal{F}:P\rightarrow\hat{\mathcal{V}},\quad\hat{\mathcal{V}}=\{\{{V}_{j},\ {A}_{j}\}\mid j=1,2,\dots,N\},$
@@
Enhance: Based on the feedback and guidelines, the generation request is refined to improve quality in the next iteration:
(4) $p_{i+1}=R(p_{i};f_{i},g),$
@@
3.2. Script Writing Guidelines
@@
Current video generation models have limited contextual modeling capacity; as a result, adjacent shots set within the same spatial environment often suffer from inconsistency in background elements. Therefore, scriptwriters should avoid placing successive shots in the same background and should minimize transitions between tightly connected spatial locations (e.g., from ”outside the door” to ”inside the room,” or from one room to another). Moreover, to reduce the jarring effect of spatial jumps, it is recommended not to place the same character in consecutive shots unless narrative necessity dictates it.
@@
Current video generation models struggle to synthesize multiple sequential actions within a single shot (e.g., running followed by walking and then dancing). Therefore, it is advisable to limit each shot to a single, simple action. Furthermore, generative models often fail to produce fine-grained visual elements accurately, such as textual content on smartphone screens or small patterns on objects. Scripts should avoid specifying such details.
@@
Within the Structure Guide and the Content Guide, a script may risk becoming fragmented or visually monotonous if not properly orchestrated. The style guide thus ensures harmonious pacing, visual diversity, and narrative logic. Moreover, the script’s tone and theme must align with the user’s request, and different genres require distinct narrative pacing—emotional and family-oriented scripts benefit from a slow, layered progression, while action-oriented scripts demand rapid and dynamic developments. Additionally, structural completeness is essential: every script must contain a title, character definitions, and a shot-by-shot script. Each shot should specify the character (or ”None” if absent), shot content, and accompanying subtitle or voice-over. Subtitles must align semantically with the shot content.
3.3. Script Writing Stage
Upon receiving a brief user prompt, four agents—the Scriptwriter, the Structure Reviewer, the Content Reviewer, and the Style Reviewer—work collaboratively in the script writing stage to produce a structured script.
@@ 3.3. Script Writing Stage
(6) $f_{\mathrm{scr}}^{i}=\left[E_{\textrm{str}}(S_{i},g_{\textrm{str}}),\ E_{\textrm{con}}(S_{i},g_{\textrm{con}}),\ E_{\textrm{sty}}(S_{i},g_{\textrm{sty}})\right],$
@@
(7) $S_{i+1}=G_{\textrm{scr}}(f_{\mathrm{scr}}^{i};\ S_{i},g_{\textrm{scr}})$
@@
3.4. Shot Designing Stage
@@
(9) $f_{\mathrm{shot}}^{i}=E_{\mathrm{shot}}(\hat{S}_{i}).$
The Shot Designer then uses this feedback to enhance the design:
(10) $\hat{S}_{i+1}=G_{\mathrm{shot}}(f_{\mathrm{shot}}^{i};\ \hat{S}_{i}).$
The process repeats until the Shot Reviewer deems the result satisfactory. The final shot design is denoted as $\hat{S}=\hat{S}_{n}$.
3.5. Character Modeling Stage
To ensure the identity consistency of characters across shots, in the character modeling stage, agents construct visual representations for each character and train a LoRA model. This process incorporates the Character Prompt Generator and Caption Generator, supported by reviewer agents for both image generation (T2I Evaluator and T2I Refiner) and video animation (I2V Evaluator and I2V Refiner).
The process begins with the Character Prompt Generator, which generates both text-to-image (T2I) and image-to-video (I2V) prompts for each character. T2I prompts aim to generate a frontal and naturally posed character image, while I2V prompts are used—together with the T2I images—to synthesize character-centric video clips that capture the subject from multiple angles. After both T2I and I2V processes, multi-view character images are sampled from the generated character video. The Caption Generator then produces structured captions for each sampled frame. These images and captions are subsequently used to train LoRA models (if needed).
Both the T2I and I2V processes follow the 3E Principle, but differ from script writing and shot designing in that each iteration produces multiple samples from the generative model pools. Specifically, in the $i$-th iteration, a set of candidate images is generated by the pool of T2I models $\mathcal{M}_{I}$ using the current T2I prompt $p_{I}^{i}$:
(11) $\mathcal{I}_{i}=\{G_{I}^{k}(p_{I}^{i}),\ G_{I}^{k}\in\mathcal{M}_{I},\ k=1,2,\dots,|\mathcal{M}_{I}|\}.$

### v4

@@
Abstract
Despite recent advances, long-sequence video generation frameworks still suffer from significant limitations: poor assistive capability, suboptimal visual quality, and limited expressiveness. To mitigate these limitations, we propose MAViS, a multi-agent collaborative framework designed to assist in long-sequence video storytelling by efficiently translating ideas into visual narratives. MAViS orchestrates specialized agents across multiple stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.In each stage, agents operate under the 3E Principle—Explore, Examine, and Enhance—to ensure the completeness of intermediate outputs. Considering the capability limitations of current generative models, we propose the Script Writing Guidelines to optimize compatibility between scripts and generative tools. Experimental results demonstrate that MAViS achieves state-of-the-art performance in assistive capability, visual quality, and video expressiveness. Its modular framework further enables scalability with diverse generative models and tools. With just a brief idea description, MAViS enables users to rapidly explore diverse visual storytelling and creative directions for sequential video generation by efficiently crafting high-quality, complete long-sequence videos. To the best of our knowledge, MAViS is the only framework that provides multimodal design output – videos with narratives and background music.
MAViS: A Multi-Agent Framework for Long-Sequence Video Storytelling
Qian Wang Virginia Tech. University USA yqwq1996@vt.edu          Ziqi Huang Nanyang Technological University Singapore ZIQI002@e.ntu.edu.sg          Ruoxi Jia Virginia Tech. University USA ruoxijia@vt.edug
Paul Debevec Netflix Eyeline Studios USA debevec@scanlinevfx.com          Ning Yuthanks: Corresponding author and project lead Netflix Eyeline Studios USA ning.yu@scanlinevfx.com
1 Introduction
Methods Agentic Script Writing Shot Designing Dynamic ID Consistency Audio Assistive
Mora Yuan et al. (2024) $\times$ $\times$ $\times$ $\times$
AesopAgent Wang et al. (2024a) $\times$ $\times$ $\times$
MovieDreamer Zhao et al. (2024) $\times$ $\times$ $\times$ $\times$ $\times$
DreamFactory Xie et al. (2024) $\times$ $\times$ $\times$
VGoT Zheng et al. (2025) $\times$ $\times$ $\times$
LCT Guo et al. (2025) $\times$ $\times$ $\times$ $\times$ $\times$
MovieAgent Wu et al. (2025b) $\times$ $\times$
@@ 1 Introduction
Table 1: Properties of MAViS vs Other Methods: We summarize seven metrics for long-sequence video storytelling works."Agentic": whether the system employs a multi-agent architecture; "Script Writing": ability generate scripts from user prompt; "Shot Designing.": support for structured shot designing and controllable narrative flow; "Dynamic": whether the generated video exhibits dynamic features; "ID Consistency": consistency in character identity across multiple shots; "Audio": whether the output includes synchronized audio; "Assistive": whether the pipeline is designed to support users in script writing, script planning, or model fine-tuning by streamlining processes.
In recent years, video generative models OpenAI (2024); MiniMax (2024); GenmoTeam (2024) have made substantial strides in text-to-video (T2V) and image-to-video (I2V) generation. While these methods demonstrate promising capabilities for shot-level creation, the generation of long videos remains a formidable challenge. Besides, the current scale of computing and data is insufficient to train a single end-to-end model capable of crafting high-quality, minute-long videos.
@@
To overcome this limitation, prior works have explored extending existing video clips KlingAI (2024); Henschel et al. (2024); Ge et al. (2022), or composing sequences of keyframes Liu et al. (2024); Yang et al. (2024b); Rahman et al. (2023); Wang et al. (2024a), or integrating storytelling and video generation Yin et al. (2023b); Zheng et al. (2025). However, these approaches suffer from repetitive motion, lack of narrative cohesion, or result in static image series with limited visual expressions. To better align narrative planning with video animation, MovieDreamer Zhao et al. (2024) adopts an auto-regressive approach to generate key frames from input scripts, while Dreamfactory Xie et al. (2024) and MovieAgent Wu et al. (2025b) propose a hierarchical script–scene–shot pipeline. However, as shown in Table 1, these methods require substantial laborious effort from users for script writing and training LoRA models to maintain ID consistency, which significantly limits their assistive capability and hinders large-scale deployment.
@@
In summary, current long-sequence video storytelling methods face three major limitations: 1) poor assistive capability, not supporting script writing, shot designing, and LoRA training to maintain ID consistency; 2) suboptimal visual quality, manifested in visual distortions, unnatural actions, and inaccurate object proportions; and 3) limited expressiveness, often resulting in repetitive actions and poorly structured shot compositions.
@@
To address this, we introduce MAViS, a multi-agent collaborative framework designed to assist in long-sequence video storytelling by efficiently translating ideas into visual narratives. MAViS encompasses various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation. Given a brief idea description from the user, the script writing stage first generates a structured script with character definitions and shot outlines. The shot designing stage defines a series of shot-level elements to guide subsequent content generation, while character modeling ensures ID consistency via LoRA-trained appearances. During the keyframe generation stage, T2I models are employed to create the initial static frames for each shot, i.e., the keyframes, which are then expanded into full video sequences in the video animation stage through I2V models, maintaining visual coherence and narrative fluency. Finally, voice-overs and background music (BGM) are added to each shot during the voice generation stage, resulting in a cohesive, long-sequence video.
@@
Crucially, all stages operate under the 3E Principle: - Explore: Initial generation of candidate outputs; - Examine: Reviewing the quality and completeness of the output; - Enhance: Iterative refinement and optimization. After multiple iterations, each stage passes the most complete version of its output to the next, enabling gradual optimization throughout the pipeline.
@@
To mitigate the limitations of current generative models in maintaining background consistency, executing complex actions, and rendering fine-grained details, we propose the Script Writing Guidelines aimed at enhancing the compatibility between scripts and generative models, ultimately improving the expressiveness of the video storytelling. As illustrated in Figure LABEL:fig:teaser, with just a brief idea description, MAViS enables users to rapidly explore diverse visual storytelling and creative directions for sequential video generation by efficiently crafting high-quality, complete long-sequence videos. Furthermore, its modular architecture ensures scalability with various generative tools and models.
@@
[…]
@@
$\bullet$ Experiments demonstrate that MAViS achieves state-of-the-art performance and wins the most user preferences in terms of expressiveness and visual quality. With just a brief idea description from the user, MAViS efficiently generates high-quality and complete long-sequence video storytelling. To the best of our knowledge, MAViS is the only framework that provides multimodal design output.
2 Related Works
Figure 2: Illustration of MAViS framework.
2.1 Video Generation
@@
Long video generation He et al. (2022); Ge et al. (2022); Wang et al. (2024c) aims to produce longer videos than previous text-to-video (T2V) ModelScopeTeam (2023); Wang et al. (2024b) and image-to-video (I2V) Girdhar et al. (2023); Xing et al. (2025); Zhang et al. (2023) models. Auto-regressive approaches Weng et al. (2024); Henschel et al. (2024) and video storytelling methods Yin et al. (2023b); Zheng et al. (2025) suffer from repetitive motions with limited narrative development and poor visual quality. Meanwhile, commercial models KlingAI (2024); Runway (2024a, b) support temporal extension. However, the duration of the generated videos remains constrained. In this work, MAViS uses cutting-edge video generation models to synthesize high-quality video clips and assembles them into long-sequence video storytelling.
2.2 AI Agents
@@
AI agents Yang et al. (2023a); Yin et al. (2023a); Li et al. (2023); Hong et al. (2023); Park et al. (2023), empowered by large language models Brown et al. (2020); Touvron et al. (2023a, b), have shown broad applicability in reasoning, planning, web interaction, and visual generation. In video generation, systems Yang et al. (2024a); Yuan et al. (2024); Wang et al. (2024a) explore world modeling, multi-agent coordination, and retrieval-augmented synthesis, yet often yield static content with limited transitions. Hierarchical pipelines Xie et al. (2024); Wu et al. (2025b) introduce script-to-shot workflows but still rely on laborious scripting and LoRA tuning. In contrast, our MAViS leverages multi-agent collaboration to enable long-sequence video storytelling.
3 MAViS
@@
We propose MAViS, a multi-agent collaborative framework for long-sequence video storytelling that consists of various stages, including script writing, shot designing, character modeling, keyframe generation, video animation, and audio generation.
@@
Given a user prompt $P$, as illustrated in Figure 2 MAViS aims to generate a long-sequence video storytelling that comprises $N$ shots:
$\mathcal{F}:P\rightarrow\hat{\mathcal{V}},\quad\hat{\mathcal{V}}=\{\{{V}_{j},\ {A}_{j}\}\mid j=1,2,\dots,N\},$ (1)
@@
Enhance: Based on the feedback and guidelines, the request is refined to improve quality in the next iteration:
$p_{i+1}=R(p_{i};f_{i},g),$ (4)
@@
3.2 Script Writing Guidelines
@@
Current video generation models have limited contextual modeling capacity; as a result, adjacent shots set within the same spatial environment often suffer from inconsistency in background elements. Therefore, scriptwriters should avoid placing successive shots in the same background and should minimize transitions between tightly connected spatial locations (e.g., from "outside the door" to "inside the room," or from one room to another). Moreover, to reduce the jarring effect of spatial jumps, it is recommended not to place the same character in consecutive shots unless narrative necessity dictates it.
@@
Current video generation models struggle to synthesize multiple sequential actions within a single shot (e.g., running followed by walking and then dancing). Therefore, it is advisable to limit each shot to a single, simple action. Furthermore, generative models often fail to craft fine-grained visual elements accurately, such as textual content on smartphone screens or small patterns on objects. Scripts should avoid specifying such details.
@@
Within the Structure Guide and the Content Guide, a script may risk becoming fragmented or visually monotonous if not properly orchestrated. The style guide thus ensures harmonious pacing, visual diversity, and narrative logic. Moreover, the script’s tone and theme must align with the user’s request, and different genres require distinct narrative pacing—emotional and family-oriented scripts benefit from a slow, layered progression, while action-oriented scripts demand rapid and dynamic developments. Additionally, structural completeness is essential: every script must contain a title, character definitions, and a shot-by-shot script. Each shot should specify the character (or "None" if absent), shot content, and accompanying subtitle or voice-over. Subtitles must align semantically with the shot content.
3.3 Script Writing Stage
Upon receiving a brief user prompt, four agents—the Scriptwriter, the Structure Reviewer, the Content Reviewer, and the Style Reviewer—work collaboratively in the script writing stage to form a structured script.
@@ 3.3 Script Writing Stage
$f_{\mathrm{scr}}^{i}=\left[E_{\textrm{str}}(S_{i},g_{\textrm{str}}),\ E_{\textrm{con}}(S_{i},g_{\textrm{con}}),\ E_{\textrm{sty}}(S_{i},g_{\textrm{sty}})\right],$ (6)
@@
$S_{i+1}=G_{\textrm{scr}}(f_{\mathrm{scr}}^{i};\ S_{i},g_{\textrm{scr}})$ (7)
@@
3.4 Shot Designing Stage
@@
$f_{\mathrm{shot}}^{i}=E_{\mathrm{shot}}(\hat{S}_{i}).$ (9)
The Shot Designer then uses this feedback to enhance the design, until the Shot Reviewer deems the result satisfactory:
$\hat{S}_{i+1}=G_{\mathrm{shot}}(f_{\mathrm{shot}}^{i};\ \hat{S}_{i}),\quad\hat{S}=\hat{S}_{n}.$ (10)
3.5 Character Modeling Stage
To ensure the ID consistency of characters across shots, in the character modeling stage, agents construct visual representations for each character and train a LoRA model. This process incorporates the Character Prompt Generator and Caption Generator, supported by reviewer agents for both image generation (T2I Evaluator and T2I Refiner) and video animation (I2V Evaluator and I2V Refiner).
The process begins with the Character Prompt Generator, which generates both text-to-image (T2I) and image-to-video (I2V) prompts for each character. T2I prompts aim to generate a frontal and naturally posed character image, while I2V prompts are used—together with the T2I images—to synthesize character-centric video clips that capture the subject from multiple angles. After both T2I and I2V processes, multi-view character images are sampled from the generated character video. The Caption Generator then generates structured captions for each sampled frame. These images and captions are subsequently used to train LoRA models (if needed).
Both the T2I and I2V processes follow the 3E Principle, but differ from script writing and shot designing in that each iteration generates multiple samples from the generative model pools. Specifically, in the $i$-th iteration, a set of candidate images is generated by the pool of T2I models $\mathcal{M}_{I}$ using the current T2I prompt $p_{I}^{i}$:
$\mathcal{I}_{i}=\{G_{I}^{k}(p_{I}^{i}),\ G_{I}^{k}\in\mathcal{M}_{I},\ k=1,2,\dots,|\mathcal{M}_{I}|\}.$ (11)